OpenStudy (anonymous):

A rectangle has an area of 12 cm^2 and a width of 400 cm. what is the length of the rectangle?

OpenStudy (anonymous):

@Babynini

OpenStudy (babynini):

Uh, how can the side be larger than the area?

OpenStudy (anonymous):

lol no clue

OpenStudy (babynini):

xD I really don't know haha

OpenStudy (babynini):

ooh! the length can be a decimal!

OpenStudy (babynini):

Let's see, try 12/400 = ?

OpenStudy (anonymous):

its 0.03

OpenStudy (anonymous):

my options are 3 cm 30 cm 300 cm 3000 cm

OpenStudy (babynini):

online homework?

OpenStudy (babynini):

All of those options are too large..

OpenStudy (anonymous):

nah lol paper hw i got 60 questions to answer

OpenStudy (anonymous):

Are you sure the options aren't in meters? If you convert to meters, its 3

OpenStudy (anonymous):

Are you sure it wasn't 12 m^2? Cause it would make sense then.

OpenStudy (anonymous):

oh yeah its that sorry i typed it wrong

OpenStudy (babynini):

hahah gwow. What's the correct problem?

OpenStudy (anonymous):

A rectangle has an area of 12m^2 and a width of 400 cm. what is the length of the rectangle?

OpenStudy (babynini):

the length will be 300 :)

OpenStudy (anonymous):

Then it would be 300

OpenStudy (anonymous):

thank u very much guys now i need to see how u did it

OpenStudy (anonymous):

i have to write one or 2 complete sentences on how i got it :(

OpenStudy (babynini):

Oh wow o.0

OpenStudy (anonymous):

i know right :( so how did u get 300

OpenStudy (anonymous):

Well, area of a rectangle is l*w=A. Since we know the A of 12 and the w of 4, we just divide \[\frac{ 12 }{ 4 }m=3m\] Since that would be 3 meters, we convert it back to centimeters which makes it 300 cm

OpenStudy (anonymous):

Thank you very much :)

OpenStudy (anonymous):

haha, no problem :)